OpenStudy (anonymous):

In a triangle ABC segment DE parallel to side AC. (Endpoints like on AB and BC respectively.) What is the length of AD if AB = 16 cm, AC = 20cm and DE = 15cm?

Directrix (directrix):

Triangle BDE ~ Triangle BAC by the AA Postulate. Lengths of corresponding sides of similar triangle are in proportion. BD/BA = DE/AC BD/16 = 15/20 BD/16 = 3/4 Cross multiply and solve for BD. Then, subtract BD from AB (16) to get the length of AD. @victorlau123